Calculate the speed of light in glycerol if the index of refraction for glycerol is 1.470.



1
Expert's answer
2021-01-27T15:34:29-0500

"v=\frac{c}{n}=\frac{299 792 458}{1.470}=203940448(m/s)" .Answer